OpenStudy (samsan9):

Solve tan 60 using the double angle theorem?

OpenStudy (anonymous):

why? you know both \(\sin(60)=\frac{\sqrt3}{2}\) and \(\cos(60)=\frac{1}{2}\) so divide the first by the second

OpenStudy (samsan9):

i have to use this exact formula for my final coming up :/